OpenStudy (anonymous):

The sum of two consecutive integers is 37. Write an equation that models this situation and find the values of the two integers.

OpenStudy (anonymous):

cheeseburgers

OpenStudy (anonymous):

consecutive means one number right after the other one. So lets let one number be x and the other number be x + 1. x + x + 1 = 37 -- combine like terms 2x + 1 = 37 -- subtract 1 2x = 36 -- divide by 2 x = 18 x + 1 = 18 + 1 = 19 check.. 18 + 19 = 36 36 = 36 (correct) your 2 numbers are 18 and 19

OpenStudy (anonymous):

cheeseburgers :)